# Inequality Restricted Minimum Density Power Divergence Estimation in Panel Count Data

### v1

@@ Acknowledgements
$-\frac{\gamma}{m}\sum_{i=1}^{m}\bm{u}(\bm{N_{i}})\bm{u}^{T}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}})+\frac{1}{m}\sum_{i=1}^{m}\bm{I}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}})$
@@
$\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{\gamma+1}(\bm{n})=\bm{J}_{\gamma}(\bm{\theta_{0}})$ (A.24)
@@
$\mathrm{E}\left[\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\right]=\mathbf{0}_{p},\quad\text{ and }\operatorname{Var}\left[\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\right]=\bm{K}_{\gamma}\left(\bm{\theta}_{0}\right)$
@@
$\bm{K}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{2\gamma+1}(\bm{n})-\bm{\xi}_{\gamma}(\bm{\theta})\bm{\xi}_{\gamma}^{T}(\bm{\theta}).\$
@@
Let us consider the following estimating equations satisfied by $\bm{\theta}$
$\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}+\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+\frac{\bm{\hat{\lambda}^{(s)}}}{\sqrt{m}}\bm{H}_{s}\left(\bm{\theta}_{0}\right)+o_{p}(1)=\mathbf{0}_{p}$ (A.26)
$\bm{H^{T}_{s;\bm{\theta}_{0}}}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+o_{p}(1)=\mathbf{0}_{s}\$ (A.27)
$\bm{H^{T}_{s^{*};\bm{\theta}_{0}}}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+o_{p}(1)=\mathbf{0}_{q}\$ (A.28)
Now we can express equations (A.26), (A.27) and (A.28) in the matrix form as
$\left(\begin{array}[]{ccc}\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)&\bm{H}_{s;\bm{\theta}_{0}}&\bm{H}_{s^{*};\bm{\theta}_{0}}\\ \bm{H}_{s;\bm{\theta}_{0}}&\mathbf{0}_{s\times s}&\mathbf{0}_{s\times q}\\ \bm{H}^{T}_{s_{*};\bm{\theta}_{0}}&\mathbf{0}_{q\times s}&\mathbf{0}_{q\times q}\end{array}\right)\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}}}{\sqrt{m}}\\ 0\end{array}\right)=\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ 0\end{array}\right)+o_{p}(1)$
Therefore
$\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}}}{\sqrt{m}}\\ 0\end{array}\right)=\left(\begin{array}[]{ccc}\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)&\bm{H}_{s;\bm{\theta}_{0}}&\bm{H}_{s^{*};\bm{\theta}_{0}}\\ \bm{H}_{s;\bm{\theta}_{0}}&\mathbf{0}_{s\times s}&\mathbf{0}_{s\times q}\\ \bm{H}^{T}_{s_{*};\bm{\theta}_{0}}&\mathbf{0}_{q\times s}&\mathbf{0}_{q\times q}\end{array}\right)^{-1}\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ 0\end{array}\right)+o_{p}(1)$
This gives
$\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}}}{\sqrt{m}}\\ 0\end{array}\right)=\left(\begin{array}[]{ccc}\bm{A}^{-1}+\bm{A}^{-1}\bm{B}\bm{A}^{-1}&-\bm{A}^{-1}\bm{C}\bm{P}&-\bm{A}^{-1}\bm{U}\bm{S}^{-1}\\ -\bm{P}^{-1}\bm{D}\bm{S}^{-1}&\bm{P}^{-1}+\bm{P}^{-1}\bm{Q}\bm{S}^{-1}\bm{R}\bm{P}^{-1}&-\bm{P}^{-1}\bm{Q}\bm{S}^{-1}\\ -\bm{S}^{-1}\bm{V}\bm{A}^{-1}&-\bm{S}^{-1}\bm{R}\bm{P}^{-1}&\bm{S}^{-1}\end{array}\right)$$\times$
$\begin{subarray}{c}\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ 0\end{array}\right)\end{subarray}+o_{p}(1)$
where,
$\begin{array}[]{c}\bm{A}=J_{\gamma}(\bm{\theta}_{0}),\quad\bm{B}=H_{s}(\bm{\theta_{0}})P^{-1}H^{T}_{s}(\bm{\theta_{0}})+US^{-1}V,\\ \bm{C}=H_{s}(\bm{\theta_{0}})-US^{-1}R,\quad\bm{D}=H^{T}_{s}(\bm{\theta_{0}})-QS^{-1}V,\\ \bm{P}=-H_{s}(\bm{\theta_{0}})J^{-1}_{\gamma}(\bm{\theta}_{0})H^{T}_{s}(\bm{\theta_{0}}),\quad\bm{Q}=-H_{s}(\bm{\theta_{0}})J^{-1}_{\gamma}(\bm{\theta}_{0})H^{T}_{s^{*}}(\bm{\theta_{0}}),\\ \bm{R}=-H_{s^{*}}(\bm{\theta_{0}})J^{-1}_{\gamma}(\bm{\theta}_{0})H^{T}_{s}(\bm{\theta_{0}}),\quad\bm{S}=-H_{s^{*}}(\bm{\theta_{0}})J^{-1}_{\gamma}(\bm{\theta}_{0})H^{T}_{s^{*}}(\bm{\theta_{0}}),\\ \bm{U}=H_{s^{*}}(\bm{\theta_{0}})-H_{s}(\bm{\theta_{0}})P^{-1}Q,\quad\bm{V}=H^{T}_{s^{*}}(\bm{\theta_{0}})-RP^{-1}H^{T}_{s}(\bm{\theta_{0}})\\ \end{array}$
Hence,
@@
$\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)=-[\bm{A}^{-1}+\bm{A}^{-1}\bm{B}\bm{A}^{-1}]\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}+o_{p}(1)$
From equation (A.25), we can write

### v4

@@ Appendix B Proof of Theorem (b)
$-\frac{\gamma}{m}\sum_{i=1}^{m}\bm{u}(\bm{N_{i}})\bm{u}^{T}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}})+\frac{1}{m}\sum_{i=1}^{m}\bm{I}(\bm{N_{i}})f^{\gamma}(\bm{N_{i}}).$
@@
$\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{\gamma+1}(\bm{n})=\bm{J}_{\gamma}(\bm{\theta_{0}}).$ (B.2)
@@
$\mathrm{E}\left[\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\right]=\mathbf{0}_{p},\quad\text{ and }\operatorname{Var}\left[\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\right]=\bm{K}_{\gamma}\left(\bm{\theta}_{0}\right),$
@@
$\bm{K}_{\gamma}(\bm{\theta})=\sum_{\begin{subarray}{c}n_{jl}=0\\ j=1,2\\ l=1,2,\dots,k\end{subarray}}^{\infty}\bm{u}(\bm{n})\bm{u}^{T}(\bm{n})f^{2\gamma+1}(\bm{n})-\bm{\xi}_{\gamma}(\bm{\theta})\bm{\xi}_{\gamma}^{T}(\bm{\theta}).\\$
@@
$\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\xrightarrow[m\rightarrow\infty]{L}\mathcal{N}\left(\mathbf{0}_{p},\bm{K}_{\gamma}\left(\bm{\theta}_{0}\right)\right).\\$ (B.3)
Let us consider the following estimating equations satisfied by $\bm{\theta}$. Here, combining (A.6) and (A.7), we have
$\left.\sqrt{m}\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}+\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+\frac{\bm{\hat{\lambda}^{\textbf{(s)}}}}{\sqrt{m}}\bm{H}_{s;\bm{\theta}_{0}}+o_{p}(1)=\mathbf{0}_{p},\\$ (B.4)
Now, from (A.16), we derive
$\bm{H^{T}_{s;\bm{\theta}_{0}}}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+o_{p}(1)=\mathbf{0}_{s},\\$ \\ (B.5)
Similarly, using (A.17), we obtain
$\sqrt{m}\left(\bm{h^{\left(\textbf{s}^{*}\right)}}(\widetilde{\bm{\theta}}_{\gamma})-\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)\right)=\bm{H^{T}_{s^{*};\bm{\theta}_{0}}}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)+o_{p}(1).\\$ \\ (B.6)
Now, we can express equations (B.4), (B.5) and (B.6) in the matrix form as
$\left(\begin{array}[]{ccc}\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)&\bm{H}_{s;\bm{\theta}_{0}}&\bm{H}_{s^{*};\bm{\theta}_{0}}\\ \bm{H}^{T}_{s;\bm{\theta}_{0}}&\mathbf{0}_{s\times s}&\mathbf{0}_{s\times q}\\ \bm{H}^{T}_{s^{*};\bm{\theta}_{0}}&\mathbf{0}_{q\times s}&\mathbf{0}_{q\times q}\end{array}\right)\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}^{\textbf{(s)}}}}{\sqrt{m}}\\ 0\end{array}\right)$
$=\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ \sqrt{m}\left(\bm{h^{\left(\textbf{s}^{*}\right)}}(\widetilde{\bm{\theta}}_{\gamma})-\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)\right)\end{array}\right)+o_{p}(1).$
Therefore,
$\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}^{\textbf{(s)}}}}{\sqrt{m}}\\ 0\end{array}\right)=\left(\begin{array}[]{ccc}\bm{J}_{\gamma}\left(\bm{\theta}_{0}\right)&\bm{H}_{s;\bm{\theta}_{0}}&\bm{H}_{s^{*};\bm{\theta}_{0}}\\ \bm{H}^{T}_{s;\bm{\theta}_{0}}&\mathbf{0}_{s\times s}&\mathbf{0}_{s\times q}\\ \bm{H}^{T}_{s^{*};\bm{\theta}_{0}}&\mathbf{0}_{q\times s}&\mathbf{0}_{q\times q}\end{array}\right)^{-1}$ $\times$
$\begin{subarray}{c}\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ \sqrt{m}\left(\bm{h^{\left(\textbf{s}^{*}\right)}}(\widetilde{\bm{\theta}}_{\gamma})-\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)\right)\end{array}\right)\end{subarray}+o_{p}(1).$
This gives
$\left(\begin{array}[]{c}\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)\\ \frac{\bm{\hat{\lambda}^{\textbf{(s)}}}}{\sqrt{m}}\\ 0\end{array}\right)=\left(\begin{array}[]{ccc}\bm{A}^{-1}+\bm{A}^{-1}\bm{B}\bm{A}^{-1}&-\bm{A}^{-1}\bm{C}\bm{P}&-\bm{A}^{-1}\bm{U}\bm{S}^{-1}\\ -\bm{P}^{-1}\bm{D}\bm{S}^{-1}&\bm{P}^{-1}+\bm{P}^{-1}\bm{Q}\bm{S}^{-1}\bm{R}\bm{P}^{-1}&-\bm{P}^{-1}\bm{Q}\bm{S}^{-1}\\ -\bm{S}^{-1}\bm{V}\bm{A}^{-1}&-\bm{S}^{-1}\bm{R}\bm{P}^{-1}&\bm{S}^{-1}\end{array}\right)$ $\times$
$\begin{subarray}{c}\left(\begin{array}[]{c}-\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}\\ 0\\ \sqrt{m}\left(\bm{h^{\left(\textbf{s}^{*}\right)}}(\widetilde{\bm{\theta}}_{\gamma})-\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)\right)\end{array}\right)\end{subarray}+o_{p}(1),$ (B.7)
where
@@
Hence, we get
@@
$\sqrt{m}\left(\widetilde{\bm{\theta}}_{\gamma}-\bm{\theta}_{0}\right)=-[\bm{A}^{-1}+\bm{A}^{-1}\bm{B}\bm{A}^{-1}]\sqrt{m}\left.\frac{\partial}{\partial\bm{\theta}}g_{m}(\bm{\theta})\right|_{\bm{\theta}=\bm{\theta}_{0}}$
$-[\bm{A}^{-1}\bm{U}\bm{S}^{-1}]\sqrt{m}\left(\bm{h^{\left(\textbf{s}^{*}\right)}}(\widetilde{\bm{\theta}}_{\gamma})-\bm{h^{\left(\textbf{s}^{*}\right)}}\left(\bm{\theta_{0}}\right)\right)+o_{p}(1).$ (B.8)
and